OpenStudy (anonymous):

An implicit equation for the line through (2,-1) normal to the vector<-3,3> is?

OpenStudy (anonymous):

Okay, first you want to find the slope of the vector. Then you can find out the slope of the line which is normal to it.

OpenStudy (anonymous):

Is the slope -1?

OpenStudy (amistre64):

implicit .... this sounds like its not going to be in slope intercept form is it

OpenStudy (amistre64):

will it be a vector equation? r = p + v or, x= ... ; y= ... or maybe ax+by=c

OpenStudy (anonymous):

I believe implicit means the same as general which is ax + by = c.

OpenStudy (amistre64):

we can live with that :)

OpenStudy (amistre64):

yes, -1 is the slope of the given vector; do you recall the definition for a normal vector?

OpenStudy (anonymous):

A normal vector.. N dot x = N dot p?

OpenStudy (amistre64):

hmmm, i was looking more for: perpendicular

OpenStudy (amistre64):

the slope of a vector and its normal vector are 90 degree to each other; at right angles; orthogonal .... perpendicular

OpenStudy (amistre64):

if you know the slope of the given vector to be -1; what property of perp slopes gives us the slope of the normal vector?

OpenStudy (anonymous):

the negative reciprocal?

OpenStudy (amistre64):

if the given slope is m1 and the perp slope is m2 then m1 * m2 = -1 yes, the negative reciprical is a good description of it :)

OpenStudy (amistre64):

so the slope of our line is the negative reciprical of -1; which is 1 correct? and when given a point and a slope; we can utilize the point slope form of a line and simplify it into the ax+by=c form from there

OpenStudy (anonymous):

Okay, so that's what I thought.. so how can I write it in ax + by = c form?

OpenStudy (amistre64):

when we have a point, (a,b); and a slope, m; the point slope form of the line takes the form:\[y-b=m(x-a)\]we can push this around to get\[y-b=mx-ma\]\[-mx+y=-ma+b\]and since some text dont like a negative out front we might have to clean it up by multiplying thru by -1

OpenStudy (amistre64):

\[-(1)x+y=-(1)(2)+(-1)\] \[-x+y=-2-1\] \[-x+y=-3\] \[x-y=3\]

OpenStudy (anonymous):



OpenStudy (amistre64):

Raph is just showing off :)

OpenStudy (anonymous):

So he isn't doing what is supposed to be done for this problem?

OpenStudy (amistre64):

dunno, there wasnt a method asked for

OpenStudy (anonymous):

I know how to out find it, but Raphael's method almost seems like the intended method.

OpenStudy (amistre64):

lol; yeah, it is pretty sweet; almost feels like cheating

OpenStudy (anonymous):

@amistre64 lol

OpenStudy (anonymous):

@RaphaelFilgueiras What inspired the (x-2, y+1) vector?

OpenStudy (anonymous):

if the point x,y lie in line then the vector(x-2,x-(-1) will lie in line too

OpenStudy (anonymous):

@RaphaelFilgueiras Can you elaborate? How do you know?

OpenStudy (anonymous):

Hmmm I know the vector components will remain proportional to each other...

OpenStudy (anonymous):

That proportion will be the slope by design.

OpenStudy (anonymous):



OpenStudy (anonymous):

@wio get it?

OpenStudy (anonymous):

Yeah I get it.